# Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level

### v1

@@ 3.3 Speculative decoding
where $\mathrm{EAGLE\_Head}$ consists of a FC layer and a decoder layer from the original LLM; $\mathbf{e}_{k}$ represents the embedding of the candidate token $x_{k}$. After predicting the hidden state at step $t_{i}$, the step-by-step safe token generation is conducted on this predicted hidden state.
@@
where $\mathbf{h}_{0}$ denotes the hidden state of the prefill at step $t_{0}$, $x_{N}$ represents the output of LLMs. Equation (7) reveals that RDS only generates the safe response from the hidden state of prefill, without additional LLMs training nor other models introduced.
@@ 3.4 Highlights
Finally, to enhance the reasoning speed of RDS and facilitate its practical implementation, we incorporate a speculative decoding structure into the resampling process. It leverages the advantages of the step-by-step mechanism to accelerate the generation process. Experimental results demonstrate that the token generation speed of RDS is approximately $\mathbf{2.12\times\sim 3.09\times}$ faster than that of the baselines. These improvements demonstrate both the effectiveness and efficiency of RDS.
@@ 4.1 Experimental setup
We evaluate the security and helpfulness of RDS and baselines on five benchmarks. (1) Custom (Zheng et al., 2024) contains 100 harmful and 100 harmless queries, which are “How to do” query pairs synthesized from gpt-3.5-turbo. (2) AdvBench (Zou et al., 2023) contains 250 harmful queries, we select the top-100 for evaluation. (3) MaliciousInstruct (Huang et al., 2023) contains 100 harmful question prompts, such as “How to …?” and “Ideas for…?”. It is specifically designed to encompass various harmful intents (a total of 10), including sabotage, theft, defamation, etc. (4) Testset (Zheng et al., 2024) contains 100 harmless queries synthesized from gpt-3.5-turbo to evaluate the false refusals of baselines. (5) Xstest (Röttger et al., 2023) contains multiple contrastive queries; we select 100 harmful queries and 100 benign queries in contrastive groups.
Models
We deploy RDS on five opensource LLMs, namely Llama-2-chat-7B (Touvron et al., 2023), Llama-3-8b-Instruct (Dubey et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024), Vicuna-7B-v1.3 (Chiang et al., 2023), and Vicuna-13B-v1.3 (Chiang et al., 2023). For brevity, we use LLaMA2, LLaMA3, Qwen2, Vicuna-7B, and Vicuna-13B to refer to them, respectively.
Table 1: Compliance results on harmful benchmarks of baselines and RDS ($\downarrow$). The best results are in bold. We sample five responses for each query. Once a response is compliant ($s_{i}>0$) is the answer treated to be compliant. “base” represents the original LLM.
Data Defense Vicuna-7B Vicuna-13B LLaMA2 LLaMA3 Qwen2 Average
Custom base 19 6 0 0 0 5.0
safety prompt 6 0 0 0 0 1.2
Self-Remind 0 0 0 0 0 0
DRO 0 1 0 0 0 0.2
Self-Examination 0 0 0 0 0 0
SafeDecoding 0 0 0 0 0 0
RDS 0 0 0 0 0 0
AdvBench base 22 10 0 1 2 7.0
safety prompt 6 0 0 0 0 1.2
Self-Remind 0 0 0 1 0 0.2
DRO 2 0 0 0 0 0.4
Self-Examination 0 0 0 0 0 0
SafeDecoding 0 0 0 0 0 0
RDS 1 0 0 0 0 0.2
Malicious Instruct base 16 3 0 0 3 4.4
safety prompt 16 3 0 0 3 4.4
Self-Remind 0 0 0 0 0 0
DRO 3 2 0 1 2 1.2
Self-Examination 0 0 0 0 0 0
SafeDecoding 0 0 0 0 0 0
RDS 1 0 0 0 1 0.4
Xstest base 48 12 0 0 12 14.4
safety prompt 0 0 0 0 0 0
Self-Remind 0 0 0 0 0 0
DRO 4 4 0 0 0 1.6
Self-Examination 0 0 0 0 0 0
SafeDecoding 0 0 0 0 0 0
RDS 0 0 0 0 0 0
@@
We select five methods as the baselines. Prefill-based defenses contain: (1) safety prompt, which is the official safety prompt of LLaMA-2 illustrated in Appendix F. The safety prompt serves as the system prompt of LLMs. (2) Self-Remind ((Wu et al., 2023a)), which encapsulates the user’s query in a system prompt to remind LLMs to respond responsibly. (3) DRO ((Zheng et al., 2024)), which utilizes the distinguished ability at the prefill level to train the safety prompt embedding to improve the moving direction of the input. Response-based defenses contain: (4) Self-Examination ((Phute et al., 2023)), which checks the response by the LLM itself. (5) SafeDecoding (Xu et al., 2024) that matches the string of safety disclaimers and amplifies the token probabilities of safety disclaimers by training an additional expert model for LLMs.
@@
We select five samples for each query. The evaluation method still follows Section 2.2 to judge whether the response is compliant. We take the average evaluation score of all samples as the final score, which can be denoted as:
$s_{i}=\frac{1}{n}{\textstyle\sum_{j=1}^{n}}s_{ij},$ (8)
where $n=5$ is the number of response samples. For harmful queries, the threshold is set to 0. For benign queries, the threshold is set to 0.5.
@@ 4.2 Main Results
Table 1 presents the baselines, and RDS compliance results on harmful datasets. “base” represents the original model. From Table 1, we have the following inclusions.
Firstly, RDS effectively reduces compliance to harmful queries on most datasets. After applying the safety prompt, the defense ability of all LLMs further improves (i.e., vicuna-7B on Xstest). However, safety prompt does not always work (i.e., vicuna-7B on MaliciousInstruct). Compared with prefill-level defenses, RDS demonstrates further advantages (0.4 vs. 1.2 on MaliciousInstruct; 0 vs. 1.6 on Xstest). Compared with response-level defense, RDS is on par with SafeDecoding. These indicate that RDS demonstrates excellent defense ability at the decoder level.
Secondly, though trained on Custom, the classifier retains its discriminative ability on other datasets, which demonstrates the transferability of the classifier and the generalization ability of RDS.
Thirdly, larger models demonstrate better initial defense than smaller ones within the same series of LLMs. For instance, in the Vicuna series models, the 13B model surpasses the 7B model in security (6 vs.19 on Custom; 10 vs. 22 on Advenbench; 24 vs. 43 on MaliciousInstruct; 12 vs. 48 on Xstest). We speculate this is because the larger model achieves better safety alignment during training on the more extensive training data.
Table 2: Refusal results on benign benchmarks of the baselines and RDS ($\downarrow$). We sample five responses for each query. The query is treated as a refusal if half of the responses ($s<0.5$) are refused.
Data Defense Vicuna-7B Vicuna-13B LLaMA2 LLaMA3 Qwen2 Average
Custom base 0 0 4 0 0 0.8
safety prompt 0 0 5 1 0 1.2
Self-Remind 1 3 5 7 1 3.4
DRO 0 0 5 0 0 1.0
Self-Examination 0 0 100 17 0 23.4
SafeDecoding 5 6 14 3 3 6.2
RDS 0 0 3 0 0 0.6
Testset base 0 0 1 0 0 0.2
safety prompt 0 2 3 0 0 1.0
Self-Remind 3 2 1 8 1 3.0
DRO 0 0 3 0 0 0.6
Self-Examination 2 1 100 10 0 22.6
SafeDecoding 4 4 16 2 3 5.8
RDS 0 0 1 0 0 0.2
Xstest base 4 20 64 12 12 22.4
safety prompt 16 28 88 36 8 35.2
Self-Remind 52 48 96 92 24 62.4
DRO 32 72 88 36 24 46.0
Self-Examination 24 28 100 48 24 44.8
SafeDecoding 64 72 96 64 60 71.2
RDS 0 12 64 12 12 20.0
4.3 Helpfulness Analysis
Table 2 presents the refusal results on the benign datasets of baselines and RDS. The higher the refusal score, the greater the loss of helpfulness caused by defense methods.
Compared to all baselines, our model shows fewer refusal results to benign queries (0.6 vs.23.4 on Custom; 0.2 vs. 22.6 on Testset; 20.0 vs. 62.4 on Xstest). Besides, RDS can even correct the original model’s overcorrection (20.0 vs. 22.4 on Xstest). Combined with Table 1, it can be included that RDS can enhance the security of LLMs without compromising its helpfulness.
What’s more, there is an interesting observation that alignment models tend to reject benign queries (LLaMA2 vs. vicuna-7B). This phenomenon is most evident in Xstest, as it contains polysemous words. For instance, “kill” is harmful when referring to “killing a person” whereas it is benign when referring to ”kill time”. The increased refusal results of the alignment models on Xstest also reveal that current safety alignment is, to some extent, word-level alignment (20.0 of RDS vs. 46.0 of DRO). While RDS can correct this over-correction at the token level. In Appendix G, we analyze the reasons for the large number of rejections on Xstest in detail.
@@ 4.4 Efficiency Analysis
We evaluate the efficiency of baselines and RDS on the Custom (benign). We select one response for each query and collect the number of tokens generated per second in Table 3.
The inference speed of prefill-level defenses nearly matches that of the original LLMs. At the same time, response-level defenses are notably lower than the original model’s when inferencing. Due to the string matching and additional inference of the expert model, SafeDecoding demands supplementary computational resources. Moreover, SafeDecoding necessitates training the expert model for each LLM, which demands exceptionally high computational resources. Conversely, RDS employs step-by-step safe token generation without additional training of LLMs, utilizing speculative decoding to enhance the inference speed, proving significantly more efficient than other baseline methods. With the support of speculative decoding, the generation speed of RDS is 2.12x $\sim$ 3.09x faster than other baselines. We design another variant, “- w/o SD”, that removes the speculative decoding of RDS. In “- w/o SD”, The inference speed significantly drops. We include the reason for the drop that “- w/o SD” predicts the hidden state of candidate tokens by autoregressive decoding.
@@
Model vicuna-7B vicuna-13B LLaMA2 LLaMA3 Qwen2
base 41.68 31.74 42.30 38.77 34.61
safety prompt 43.43 32.15 43.09 40.20 36.43
Self-Remind 31.85 25.62 32.27 29.15 40.81
@@ 5.1 Existing Defenses
Prefill-level defenses induce LLMs to reject harmful questions by optimizing the input, such as adding a safety system prompt or filtering the input. For instance, IAPrompt ((Zhang et al., 2024b)) delves into the intent of input before decoding. Perplexity filtering ((Alon & Kamfonas, 2023)) proposes to detect the adversarial suffixes as the signal of harmful input before generating a response. Modifying the input on the encoder for defense can also be modified for attack ((Zhao et al., 2024)). Modifying input for attacks is a deficient consumption and efficient behavior. At present, multiple methods have successfully carried out jailbreak attacks from user input, such as GCG ((Zou et al., 2023)), Auto-DAN ((Zhu et al., 2023)), Evil Geniuses ((Tian et al., 2023)). Besides, input-based defenses show poor helpfulness with over-defense ((Zhou et al., 2024)).
Response-level defenses enhance the security of LLMs by judging the generated response, which follows the paradigm of generate and then judge. For instance, Self-Examination ((Phute et al., 2023)) checks the response response itself by a pre-defined prompt. SelfDecoding ((Xu et al., 2024)) captures the safety disclaimers from top tokens and amplifies their token probabilities. Response-level defenses must fully generate the output before judging, which affects the model’s efficiency. While RDS monitors the token step-by-step, forcing safe token generation in time.
@@ 5.2 Speculative Decoding
Traditionally, token generation is performed step-by-step, where the model generates one token for each step by autoregressive decoding. The generated token concatenated to the input serves as the new input for the next step ((Chen et al., 2023a)). This approach is straightforward but can be computationally expensive and slow, particularly when generating long text ((Kim et al., 2023)).
Speculative Decoding is an optimization technique used in LLMs to accelerate the process of token generation ((Leviathan et al., 2023; Chen et al., 2023b)). By the Draft-then-Verify paradigm, speculative decoding generates multiple tokens at each step ((Xia et al., 2024)). For example, (Zhang et al., 2024a) proposes to use the same serious but more minor LLM as the draft model without additional training. Not all models have a smaller draft model; self-draft becomes a new paradigm instead of using a separate draft model. For instance, Medusa ((Cai et al., 2024)) incorporates feedforward neural heads atop the decoder to predict tokens in different positions in parallel.
@@ 6 Conclusions
Our study delves into and confirms the discriminative capacity of LLMs at the decoder level. Through preliminary validation, we indicate that LLMs consistently can discern the harmfulness of output tokens at multiple steps. Motivated by these findings, we propose a Root Defense Strategy originating from the decoding level, namely RDS. The incremental safe token generation process enforces security measures. Furthermore, speculative decoding is introduced in RDS to enhance usability and facilitate deployment. Comparative experimental results demonstrate that RDS offers robust and efficient security defense without compromising utility.
@@ Appendix A ETHICS STATEMENT
Appendix B Case study
In this section, we compare baseline and RDS from the perspective of the semantics of refusal response. As shown in Figure 4, there is compliance for the harmful query even if a safety prompt is added to the prefill. Although other baselines have rejected the query, they are semantically singular. However, RDS has started with “The best way of stealing a car is”, but in the following tokens, it gives a rejection of “not advisable”, which refuses to answer the harmful query. This reflects the step-by-step safe token generation of RDS. Furthermore, RDS gives a kind hint after the rejection, which is more in line with human common sense. The above results show the superiority of RDS on semantics.
The results of the symmetry harmless problem on Xstest are shown in Figure 5. For the same behavior “stealing a car”, it is a harmful behavior in a real-world scenario. However, LLMs serve as tools and must respond to video game user requests. We found that all defenses except RDS give a rejection. It is worth noting that although SafeDecoding recognizes GPT4 as a game, it still rejects the query. At the same time, when it replies, RDS gives a friendly hint ( i.e., “it is illegal and not endorsed by Rockstar Games”).
Appendix C Training process of the classifier
$\mathbf{u}=\frac{1}{N}{\textstyle\sum_{i=1}^{N}}\mathbf{h}_{i},$ (9)
$\mathbf{m_{i}}={\mathbf{V}}^{T}(\mathbf{h_{i}}-\mathbf{u}),$ (10)
$\hat{c_{i}}={\mathbf{W}}^{T}\mathbf{m_{i}}+\mathbf{b},$ (11)
$\mathcal{L}(c_{i},\hat{c}_{i})=\frac{1}{N}{\textstyle\sum_{i=1}^{N}}(c_{i}\log{\hat{c}_{i}}),$ (12)
where $\mathbf{u}\in\mathbb{R}^{d}$ is the mean value of all hidden states of queries, and $N$ is the number of queries. $\mathbf{V}\in\mathbb{R}^{d\times m}$ represents the $m$ principal components. $\mathbf{W}\in\mathbb{R}^{1\times d}$ and $\mathbf{b}$ correspond to the trainable parameters of the classifier. $\hat{c}_{i}$ and $c_{i}$ represent the predicted score and the query label.
Figure 4: An example of baselines and RDS for a harmful query. The user inputs “What is the best way of stealing a car?” as the query. The safety prompt fails to defend against this harmful query. At the same time, other defenses offer refusal answers.

### v2

@@ 3.3 Hidden State Prediction
where $\mathrm{EAGLE\_Head}$ consists of a fully-connected layer and a decoder layer from the original LLM; $\mathbf{e}_{k}$ is the embedding of the candidate token $x_{k}$. After predicting the hidden state at step $t_{i}$, the step-by-step safe token generation is conducted on this predicted hidden state.
@@
where $\mathbf{h}_{0}$ denotes the hidden state of the prefill at step $t_{0}$, $x_{N}$ represents the output of LLMs. Equation (11) reveals that RDS only generates the safe output from the hidden state of prefill, without additional LLMs training nor other models introduced.
@@ 3.4 Highlights
Finally, to enhance the reasoning speed of RDS and facilitate its practical implementation, we incorporate a speculative head into the prediction of hidden states of the candidate tokens. It leverages the advantages of the step-by-step mechanism to accelerate the generation process. Experimental results demonstrate that the token generation speed of RDS is approximately $\mathbf{2.12\times\sim 3.09\times}$ faster than that of the baselines, which demonstrates both the effectiveness and efficiency of RDS.
@@ 4.1 Experimental setup
We evaluate the security improved by different defense strategies on three harmful benchmarks: HEx-PHI (Qi et al., 2023), AdvBench (Zou et al., 2023), MaliciousInstruct (Huang et al., 2023). We assess the impact of LLMs after applying defense methods on two benign datasets: Held-out (Zheng et al., 2024), Xstest (Röttger et al., 2023). In addition, we evaluate the helpfulness of the output on Just-Eval (Lin et al., 2023) from the aspects of helpfulness, clarity, factuality, depth, and engagement.
@@
We select five defense methods as the baselines. Prefill-based defenses contain: (1) safety prompt, which is the official safety prompt of LLaMA-2 illustrated in Appendix E. The safety prompt serves as the system prompt of LLMs. (2) Self-Reminder (Wu et al., 2023a), which encapsulates the user’s query in a system prompt to remind LLMs to respond responsibly. (3) DRO (Zheng et al., 2024), which utilizes the distinguished ability at the prefill level to train the safety prompt embedding to improve the moving direction of the input. Output-based defenses contain: (4) Self-Examination (Phute et al., 2023), which checks the output by the LLM itself and filter out harmful output. (5) SafeDecoding (Xu et al., 2024), which amplifies the sampling probabilities of the output that matches the string of safety disclaimers learned from an additional trained export model.
@@
In the main results, we select 5 samples for each query and follow the evaluation strategy in Section 2.2 to judge whether a output is compliant. For Just-Eval, we use the official prompt and GPT-4 as the evaluator to score the output from 1 to 5 in terms of helpfulness, clarity, factuality, depth, and engagement.
@@ 4.2 Main Results
Table 1 presents the compliance ratio on harmful benchmarks and refusal ratio on benign benchmarks of the baselines and RDS. From Table 1, we have the following inclusions.
Firstly, RDS demonstrates excellent defense ability at the decoder level. Compared with other baselines, RDS effectively reduces compliance to harmful queries, particularly with regard to LLMs that exhibit suboptimal initial performance (i.e., Vicuna-7B). Safety prompt does not always work (i.e., Vicuna-7B on MaliciousInstruct). Furthermore, baselines reliant on the LLMs’ self-assessment, such as DRO, exhibit varying degrees of performance degradation due to the subpar capabilities of LLM itself. While RDS leverages the discriminative capabilities at the decoding level for security defense, regardless of the functionality of LLM itself. Though trained on Custom, the classifier still works on out-of-domain datasets, which demonstrates the transferability of the classifier and the generalization of RDS.
Secondly, RDS conducts security defense without increasing the rejection rate. SafeDecoding will select the matched rejection output and ignore whether the query is harmful or not. Therefore, SafeDecoding tends to reject benign query. DRO/Self-Examination relies on the initial classification ability of LLMs on input/output. In contrast, RDS solely assesses the outputs for defense, disregarding inputs. Thus, the superior utility of RDS underscores the benefits of defense mechanisms at the decoder level.
Table 1: Evaluation results on harmful and benign benchmarks. We report the percentages of harmful/benign queries where models generate compliance/refusal outputs in 5 samplings.
Model Defense Compliance on Harmful Queries ($\downarrow$) Refusal on Harmless Queries ($\downarrow$)
HEx-PHI Advbench
Malicious
Instruct
Average Held-out   Xstest Average
No defense 89 22 16 42.3 0 4 2.0
safety prompt 37 6 16 19.7 0 16 8.0
Self-Reminder 41 0 0 13.7 3 52 27.5
Vicuna-7B DRO 33 2 3 12.7 0 32 16.0
Self-Examination 23 0 0 7.7 2 24 13.0
SafeDecoding 21 0 0 7.0 4 64 24.0
RDS 16 0 0 5.3 0 0 0
No defense 46 22 16 28.0 0 20 10.0
safety prompt 14 6 16 12.0 2 28 15.0
Self-Reminder 11 0 0 3.7 2 48 25.0
Vicuna-13B DRO 3 2 3 2.7 0 72 36.0
Self-Examination 5 0 0 1.7 1 28 14.5
SafeDecoding 6 0 0 2.0 4 72 38.0
RDS 4 0 0 1.3 0 12 6.0
No defense 13 2 3 6.0 0 12 6.0
safety prompt 0 0 3 1.0 0 8 4.0
Self-Reminder 0 0 0 0 1 24 12.5
Qwen2 DRO 0 0 2 0.6 0 24 12.0
Self-Examination 0 0 0 0 0 24 12.0
SafeDecoding 0 0 0 0 3 60 31.5
RDS 0 0 0 0 0 12 6.0
No defense 27 0 0 9.0 1 64 32.5
safety prompt 0 0 0 0 3 88 45.5
Self-Reminder 0 0 0 0 1 96 48.5
Llama2 DRO 13 0 0 4.3 3 88 45.5
Self-Examination 0 0 0 0 100 100 100.0
SafeDecoding 0 0 0 0 16 96 56.0
RDS 0 0 0 0 1 64 32.5
No defense 5 1 0 2.0 0 12 6.0
safety prompt 0 0 0 0 0 36 18.0
Self-Reminder 0 1 0 0.3 8 92 50.0
Llama3 DRO 0 0 1 0.3 0 36 18.0
Self-Examination 0 0 0 0 10 48 29.0
SafeDecoding 0 0 0 0 2 64 33.0
RDS 0 0 0 0 0 12 6.0
Table 2: Evaluation results on Just-Eval. We analyze the output for benign queries from the aspect of helpfulness, clarity, factuality, depth, and engagement.
Model Defense Helpfulness Clarity Factuality Depth Engagement Average
Vicuna-13B No defense 4.55 4.87 4.48 4.28 4.29 4.49
DRO 3.90 4.69 4.12 3.37 3.89 3.99
Self-Examination 4.58 4.87 4.46 4.34 4.26 4.50
SafeDecoding 4.23 4.87 4.35 4.00 4.18 4.33
RDS 4.41 4.78 4.36 4.16 4.20 4.38
Llama2 No defense 4.59 4.95 4.42 4.51 4.67 4.63
DRO 3.52 4.59 4.00 3.06 4.13 3.86
Self-Examination 1.35 3.53 2.50 1.32 1.62 2.06
SafeDecoding 4.59 4.92 4.36 4.58 4.51 4.59
RDS 4.24 4.83 4.30 4.16 4.57 4.42
4.3 Utility Analysis
Table 2 evaluates the impact of security defense methods applied to Vicuna-13B and Llama2 across five aspects, respectively. We observe that the utility of LLMs remains largely unaffected after applying RDS. In contrast, the application of other security defense methods has led to performance degradation in LLMs to some extent, most notably in Llama2. We attribute the degradation to the over-sensitivity of these defenses, which aligns with the high refusal on harmless queries as shown in Table 1. These findings demonstrate that the resampling strategy of RDS does not compromise the semantics of the output and preserves the utility of LLMs during safety defense.
@@ 4.4 Efficiency Analysis
We evaluate the efficiency of RDS and report the number of tokens generated per second in Table 3. With the support of speculative decoding, the generation speed of RDS is 2.12x $\sim$ 3.09x faster than other baselines. We design another variant, “- w/o SD”, that removes the speculative decoding of RDS. In “- w/o SD”, The inference speed significantly drops. We include the reason for the drop that “- w/o SD” predicts the hidden state of candidate tokens by autoregressive decoding.
@@
Defense Vicuna-7B Vicuna-13B LLaMA2 LLaMA3 Qwen2
No defense 41.68 31.74 42.30 38.77 34.61
Self-Reminder 31.85 25.62 32.27 29.15 40.81
@@
4.5 Case study
Figure 4 showcases the outputs of defense methods for an example harmful query. Even safety prompt has been added to the prompt, LLMs fail to reject this harmful. Other safety defenses are semantically singular to reject the harmful query with the same rejection template. Though starting with “The best way of stealing a car is”, RDS gives a rejection of “not advisable” in the following tokens to the harmful query. This reflects that RDS identifies the harmful output during the inference and corrects it to safe tokens step-by-step.
Figure 4: Case study of baselines and RDS for a harmful query: “What’s the best way of stealing a car?”.
@@ 5.1 Existing Defenses
Prefill-level defenses induce LLMs to reject harmful questions by optimizing the input, such as adding a safety system prompt or filtering the input. For instance, IAPrompt (Zhang et al., 2024b) delves into the intent of input before decoding. Perplexity filtering (Alon and Kamfonas, 2023) proposes to detect the adversarial suffixes as the signal of harmful input before generating a output. However, prefill-level defenses can be broken through by prefill-level attack (Zhao et al., 2024). At present, multiple methods have successfully carried out jailbreak attacks from user input, such as GCG (Zou et al., 2023), Auto-DAN (Zhu et al., 2023), Evil Geniuses (Tian et al., 2023). Besides, input-based defenses show poor helpfulness with over-defense (Zhou et al., 2024).
Output-level defenses enhance the security of LLMs by judging the generated output, which follows the paradigm of generate then judge. For instance, Self-Examination (Phute et al., 2023) checks the output itself by a pre-defined prompt. SafeDecoding (Xu et al., 2024) captures the safety disclaimers and amplifies their sampling probabilities. Output-level defenses must fully generate the output before judging, which affects the model’s efficiency. While RDS monitors the token step-by-step, forcing safe token generation in time.
@@ 5.2 Speculative Decoding
Traditionally, token generation is performed step-by-step, where the model generates one token for each step by autoregressive decoding. The generated token concatenated to the input serves as the new input for the next step (Chen et al., 2023a). This approach is straightforward but can be computationally expensive and slow, particularly when generating long text (Kim et al., 2023).
Speculative Decoding is an optimization technique used in LLMs to accelerate the process of token generation (Leviathan et al., 2023; Chen et al., 2023b). By the Draft-then-Verify paradigm, speculative decoding generates multiple tokens at each step (Xia et al., 2024). For example, Tinyllama (Zhang et al., 2024a) proposes to use the same serious but more minor LLM as the draft model without additional training. Not all models have a smaller draft model; self-draft becomes a new paradigm instead of using a separate draft model. For instance, Medusa (Cai et al., 2024) incorporates feedforward neural heads atop the decoder to predict tokens in different positions in parallel.
@@ 6 Conclusions
Our study delves into and confirms the discriminative capacity of LLMs at the decoder level. Through preliminary validation, we indicate that LLMs consistently can discern the harmfulness of output tokens at multiple steps. Motivated by these findings, we propose a Root Defense Strategy originating from the decoding level, namely RDS. The incremental safe token generation process enforces security measures. Furthermore, speculative decoding is introduced in RDS to enhance usability and facilitate deployment. Comparative experiments demonstrate that RDS offers robust and efficient security defense without compromising utility.
7 Limitations
@@ Appendix A Ethics statement
Appendix B Case study on Xstest
The results of the symmetry benign problem on Xstest are shown in Figure 5. For the same behavior “stealing a car”, it is a harmful behavior in a real-world scenario. However, LLMs serve as tools and must respond to video game user requests. We found that all defenses except RDS give a rejection. It is worth noting that although SafeDecoding recognizes GPT4 as a game, it still rejects the query. At the same time, when it replies, RDS gives a friendly hint ( i.e., “it is illegal and not endorsed by Rockstar Games”).